# Heat propagation in rotating relativistic bodies

### v1

@@ I Introduction
Given the aforementioned challenges, state-of-the-art simulations of heat transport in neutron stars Schaab and Weigel (1998); Negreiros et al. (2012); Potekhin et al. (2015); Beznogov et al. (2023) employ the following workaround, originally proposed in Miralles et al. (1993). First, express the equations of the underlying microscopic theory (in their case, radiation kinetic theory) in a coordinate system tailored to the symmetries of the problem. Next, consider the energy conservation equation, and apply a series of simplifications and approximations, which invariably involve discarding certain time derivatives. The final result is an equation that resembles a diffusion equation in the chosen coordinates. Such an equation lacks general covariance and causality, forcing the user to stick to the selected coordinates, with no possibility to, e.g., boost (as the boosted initial value problem would be ill-posed). Furthermore, this equation is not guaranteed to be universal, since its derivation relies on a specific microscopic model.
@@ II.1 The relevant conservation law of the problem
Our starting point is the rigidity assumption, according to which the vector field ${\mathcal{K}}^{\mu}={\mathcal{K}}u^{\mu}$ obeys the Killing equation $\nabla_{(\mu}{\mathcal{K}}_{\nu)}=0$.111 Note that requiring ${\mathcal{K}}u^{\mu}$ to be a Killing vector also implies that the body’s rotation rate must remain constant Herglotz (1910); Noether (1910). In fact, in special relativity, any change in the rotation rate leads to a variation in the ratio “perimeter”/“radius” (due to length contraction), thereby violating the rigidity assumption. Consequently, equation (3) is valid only in systems where the temperature of the medium is much lower than the mass of its constituent particles. If this condition is not met, heat exchange can significantly alter the body’s moment of inertia (since $E=mc^{2}$”), enforcing a change in rotation rate to conserve angular momentum. This would violate the condition that ${\mathcal{K}}u^{\mu}$ is a Killing vector, thereby breaking Born rigidity. Another reason for assuming low temperature relative to mass is that it allows us to neglect the gravitational field variations induced by heat transfer. Otherwise, such variations would likely eliminate any exact spacetime symmetries, and no Killing vectors would exist. Combining this with energy conservation, we find that the four-current $J_{H}^{\mu}=-T^{\mu\nu}{\mathcal{K}}_{\nu}$ is conserved (Hawking and Ellis, 2011, §3.2), namely
@@
We emphasize that the rigidity assumption was crucial for deriving a conserved energy current whose evolution is governed only by heat transfers. In a general flow, the divergence of $\varepsilon u^{\mu}\,{+}\,q^{\mu}$ is also sensitive to the value of $P^{\mu\nu}$, which reflects the fact that thermal energy can be produced via “PdVPdV” work or via viscous heating (Landau and Lifshitz, 2013b, §50).
@@ II.2 Equilibrium states
Any theory of heat conduction that is consistent with the Tolman-Ehrenfest effect must admit (7) as a solution. Additionally, assumption (c) of Theorem 1 states that, along the solution (7), the heat flux $q^{\mu}$ should vanish, giving
@@ III Application to relativistic stars
As a consistency check, let us verify that, when we specialize (3) to relativistic stars, we obtain a hyperbolic generalization of the standard equations that are used in the neutron-star literature Schaab and Weigel (1998); Negreiros et al. (2012); Potekhin et al. (2015); Beznogov et al. (2023); Miralles et al. (1993).
@@ III.1 Non-rotating case
which is the usual equation used in neutron-star models Potekhin et al. (2015); Lander and Andersson (2018); Pons and Viganò (2019), except for an additional term with two time derivatives. Such a new term is needed for covariance, because it keeps the equations hyperbolic and causal, which guarantees that the initial value problem remains well-posed if we decide to change coordinates. On the other hand, it should be noted that, if $T$ solves (24), then $\partial_{t}(e^{\Phi}T)=\mathcal{O}(\partial^{2})$. Hence, the quantity $\partial_{t}\big{[}\kappa\partial_{t}(e^{\Phi}T)\big{]}$ really scales like a third-order derivative term (as long as $T$ is a solution of the field equation). Therefore, if we stick to the coordinates $\{t,r,\theta,\phi\}$, we may reabsorb the last term into the error $\mathcal{O}(\partial^{3})$, and we finally obtain the usual parabolic equation
@@ IV.1 Thermal evolution equation
according to which the temperature increases with distance from the rotation axis (Cercignani and Kremer, 2002, §12.5.a). This phenomenon is known as the “inertial of heat” Israel (1981). Thus, applying Theorem 1, we obtain the following evolution equation:
@@ IV.2 First application: Stationary heat flux across a cylindrical shell
Figure 1: Temperature (left panel) and heat flux (right panel) across a cylindrical shell rotating with outer tangential speed $\Omega\rho_{\text{out}}=0$ (dashed), 0.7 (blue), 0.9 (magenta), 0.99 (red), and 0.99999 (orange). For the sake of illustration, the boundary conditions have been fixed as follows: $T_{\text{out}}/T_{\text{in}}=0.5$, and $\varrho_{\text{in}}/\varrho_{\text{out}}=0.1$.
@@ IV.3 Second application: Trend to equilibrium
Figure 2: Left panel: Stationary temperature profile of a cylindrical shell with $\varrho_{\text{in}}/\varrho_{\text{out}}=0.1$ and $\Omega=0.999$, in thermal contact with an inner and an outer thermostat having the same $T$. Right panel: Thermal evolution of the same shell if the thermostats are quickly removed, and replaced with an adiabatic insulator. The various curves are snapshots at different times following the change in boundary conditions, namely $Dt/\varrho_{\text{out}}^{2}=0$ (blue), 0.05 (magenta), 0.1 (red), 0.2 (orange), $\infty$ (dashed).
@@ IV.4 Third application: Cooling of a hot cylinder with thermostatic boundary
Figure 3: Thermal evolution of a hot infinite cylinder whose surface is in contact with a thermostat with temperature $T_{\text{ther}}$. The dashed line is the non-rotating case, while the blue line refers to a cylinder spinning with surface speed $\Omega\varrho_{\text{cyl}}=0.999$.
@@ V.1 Cooling formula in curved spacetime
This is the cooling law of rigidly rotating black bodies in general relativity, expressed in a manifestly covariant form.
@@ VI Conclusions
A longstanding open problem in relativistic thermodynamics has been resolved: the derivation of a first-principles model for heat propagation in rotating rigid media. We have obtained a fully covariant equation of motion for the temperature, which generalizes and improves upon the non-covariant cooling models currently employed in neutron star physics. This equation is hyperbolic, causal, and stable, and it admits a well-posed initial-value formulation for initial data on arbitrary Cauchy surfaces. Moreover, we have demonstrated that any alternative theory of heat transport must reduce to our formulation at leading order in the derivative expansion. In other words, equation (3) cannot be further improved, unless one is willing to include additional degrees of freedom or go to higher orders in the derivative expansion. These results were only made possible by recent advances in our global understanding of relativistic fluid dynamics Bemfica et al. (2022); Kovtun (2019); Gavassino (2022a); Başar et al. (2024); Gavassino (2025).
@@ Acknowledgements
This work is partially supported by a Vanderbilt Seeding Success Grant.

### v2

@@ I Introduction
Given the aforementioned challenges, state-of-the-art simulations of heat transport in neutron stars Schaab and Weigel (1998); Negreiros et al. (2012); Yakovlev and Pethick (2004); Potekhin et al. (2015); Beznogov et al. (2023) employ the following workaround, originally proposed in Miralles et al. (1993). First, express the equations of the underlying microscopic theory (in their case, radiation kinetic theory) in a coordinate system tailored to the symmetries of the problem. Next, consider the energy conservation equation, and apply a series of simplifications and approximations, which invariably involve discarding certain time derivatives. The final result is an equation that resembles a diffusion equation in the chosen coordinates. Such an equation lacks general covariance and causality, forcing the user to stick to the selected coordinates, with no possibility to, e.g., boost (as the boosted initial value problem would be ill-posed). Furthermore, this equation is not guaranteed to be universal, since its derivation relies on a specific microscopic model.
@@
We note that the mathematical assumptions underlying Theorem 1 implicitly rely upon physical assumptions that define (and put limitations on) the domain of applicability of (3). For instance, the fact that the field equation depends only on $\{u^{\mu},{\mathcal{K}},T,n\}$ implies that local anisotropies associated with lattice structures or magnetic fields can be neglected. Moreover, the requirement that ${\mathcal{K}}u^{\mu}$ be a timelike Killing vector indicates that the metric can be treated as a stationary background, whose geometry is not affected by the transfers of heat.
@@ II.1 The relevant conservation law of the problem
Our starting point is the rigidity assumption, according to which the vector field ${\mathcal{K}}^{\mu}={\mathcal{K}}u^{\mu}$ obeys the Killing equation $\nabla_{(\mu}{\mathcal{K}}_{\nu)}=0$. [^1] Note that requiring ${\mathcal{K}}u^{\mu}$ to be a Killing vector also implies that the body’s rotation rate must remain constant Herglotz (1910); Noether (1910). In fact, in special relativity, any change in the rotation rate leads to a variation in the ratio “perimeter”/“radius” (due to length contraction), thereby violating the rigidity assumption. Consequently, equation (3) is valid only in systems where the temperature of the medium is much lower than the mass of its constituent particles. If this condition is not met, heat exchange can significantly alter the body’s moment of inertia (since $E=mc^{2}$”), enforcing a change in rotation rate to conserve angular momentum. This would violate the condition that ${\mathcal{K}}u^{\mu}$ is a Killing vector, thereby breaking Born rigidity. Another reason for assuming low temperature relative to mass is that it allows us to neglect the gravitational field variations induced by heat transfer. Otherwise, such variations would likely eliminate any exact spacetime symmetries, and no Killing vectors would exist. Combining this with energy conservation, we find that the four-current $J_{H}^{\mu}=-T^{\mu\nu}{\mathcal{K}}_{\nu}$ is conserved (Hawking and Ellis, 2011, §3.2), namely
@@
We emphasize that the rigidity assumption is crucial for deriving a conserved energy current whose evolution is governed solely by heat transfer. In a general flow, the divergence of $\varepsilon u^{\mu}\,{+}\,q^{\mu}$ is also sensitive to the value of $P^{\mu\nu}$, which reflects the fact that thermal energy can be produced via “PdVPdV” work or via viscous heating (Landau and Lifshitz, 2013b, §50).
@@ II.2 Equilibrium states
Any theory of heat conduction that is consistent with the Tolman-Ehrenfest effect must admit (7) as a solution. Additionally, assumption (c) of Theorem 1 states that, along the solution (7), the heat flux $q^{\mu}$ vanishes, giving
@@ III Application to relativistic stars
As a consistency check, let us verify that, when we specialize (3) to relativistic stars, we obtain a hyperbolic generalization of the standard equations that are used in the neutron-star literature Schaab and Weigel (1998); Negreiros et al. (2012); Yakovlev and Pethick (2004); Potekhin et al. (2015); Beznogov et al. (2023); Miralles et al. (1993).
@@ III.1 Non-rotating case
which is the usual equation used in neutron-star models Yakovlev and Pethick (2004); Potekhin et al. (2015); Lander and Andersson (2018); Pons and Viganò (2019), except for an additional term with two time derivatives. Such a new term is needed for covariance, because it keeps the equations hyperbolic and causal, which guarantees that the initial value problem remains well-posed if we decide to change coordinates. On the other hand, it should be noted that, if $T$ solves (24), then $\partial_{t}(e^{\Phi}T)=\mathcal{O}(\partial^{2})$. Hence, the quantity $\partial_{t}\big[\kappa\partial_{t}(e^{\Phi}T)\big]$ really scales like a third-order derivative term (as long as $T$ is a solution of the field equation). Therefore, if we stick to the coordinates $\{t,r,\theta,\phi\}$, we may reabsorb the last term into the error $\mathcal{O}(\partial^{3})$, and we finally obtain the usual parabolic equation
@@ IV.1 Thermal evolution equation
according to which the temperature increases with distance from the rotation axis (Cercignani and Kremer, 2002, §12.5.a). This phenomenon is known as the “inertia of heat” Israel (1981). Thus, applying Theorem 1, we obtain the following evolution equation:
@@ IV.2 First application: Stationary heat flux across a cylindrical shell
Figure 1: Temperature (left panel) and heat flux (right panel) across a cylindrical shell rotating with outer tangential speed $\Omega\varrho_{\text{out}}=0$ (dashed), 0.7 (blue), 0.9 (magenta), 0.99 (red), and 0.99999 (orange). The corresponding analytical expressions are provided in equations (36) and (37). For the sake of illustration, the boundary conditions have been fixed as follows: $T_{\text{out}}/T_{\text{in}}=0.5$, and $\varrho_{\text{in}}/\varrho_{\text{out}}=0.1$. We observe that, at high rotation rates, $T(\varrho)$ tends to form a local minimum near the outer boundary.
@@ IV.3 Second application: Trend to equilibrium
Figure 2: Left panel: Stationary temperature profile (36) of a cylindrical shell with $\varrho_{\text{in}}/\varrho_{\text{out}}=0.1$ and $\Omega=0.999$, in thermal contact with an inner and an outer thermostat having the same $T$. Right panel: Thermal evolution (40) of the same shell if the thermostats are quickly removed, and replaced with an adiabatic insulator. The various curves are snapshots at different times following the change in boundary conditions, namely $Dt/\varrho_{\text{out}}^{2}=0$ (blue), 0.05 (magenta), 0.1 (red), 0.2 (orange), $\infty$ (dashed). The temperature profile spontaneously evolves to the Tolman-Ehrenfest prediction (31), as expected.
@@ IV.4 Third application: Cooling of a hot cylinder with thermostatic boundary
Figure 3: Thermal evolution (41) of a hot infinite cylinder whose surface is in contact with a thermostat with temperature $T_{\text{ther}}$. The dashed line is the non-rotating case, while the blue line refers to a cylinder spinning with surface speed $\Omega\varrho_{\text{cyl}}=0.999$. At high rotation rates, the temperature is initially driven outward by the centrifugal term ($q^{\mu}\sim-u^{\nu}\nabla_{\nu}u^{\mu}$). Subsequently, the temperature profile develops a local minimum (as in figure 1), before relaxing to the Tolman–Ehrenfest solution (31).
@@ V.1 Cooling formula in curved spacetime
This is the cooling law of rigidly rotating black bodies in general relativity, expressed in a manifestly covariant form. The main assumptions we used to arrive here are: (a) rigid motion, (b) quasistatic evolution, (c) pure surface emission, and (d) negligible self-irradiation (and the various assumptions it entails, see footnote 3).
@@ VI Conclusions
A longstanding open problem in relativistic thermodynamics has been resolved: the derivation of a first-principles model for heat propagation in rotating rigid media. We have obtained a fully covariant equation of motion for the temperature, which generalizes and improves upon the non-covariant cooling models currently employed in neutron star physics. This equation is hyperbolic, causal, and stable, and it admits a well-posed initial-value formulation for initial data on arbitrary Cauchy surfaces. Moreover, we have demonstrated that any alternative theory of heat transport must reduce to our formulation at leading order in the derivative expansion. In other words, equation (3) cannot be further improved, unless one is willing to include additional degrees of freedom or go to higher orders in the derivative expansion. These results were only made possible by recent advances in our global understanding of relativistic fluid dynamics [^4] Our approach closely parallels that of BDNK theory Bemfica et al. (2022); Kovtun (2019), as it is formulated within a first-order derivative expansion in which the definition of temperature is fixed a posteriori to ensure causality, stability, and well-posedness. Extending the expansion to second order would lead to a Burnett-type theory for heat conduction, analogous in spirit to (non-resummed) BRSSS Baier et al. (2008), but formulated in a causal hydrodynamic frame. In contrast, our framework departs substantially from transient hydrodynamic theories (e.g. Cattaneo’s theory Cattaneo (1958), Israel–Stewart theory Israel and Stewart (1979); Hiscock and Lindblom (1983); Denicol et al. (2012), divergence-type theories Liu et al. (1986); Geroch and Lindblom (1990), and Carter’s multifluid theory Lopez-Monsalvo and Andersson (2011); Carter (1991); Priou (1991)), which treat the heat flux $q^{\mu}$ as an independent dynamical variable. Nonetheless, when such theories are truncated to first order in derivatives, they are all known to reproduce equation (20), which is equivalent to (3) up to a field redefinition. Bemfica et al. (2022); Kovtun (2019); Gavassino (2022a); Başar et al. (2024); Gavassino (2025).
@@
As an astrophysical application, equation (3) can be used to describe the cooling of self-gravitating, rapidly rotating compact objects (such as millisecond-period neutron stars) provided that the spacetime metric remains approximately unaffected by temperature changes (otherwise the existence of a timelike Killing vector field would be invalidated). This “fixed-metric approximation” is not a severe simplification, as it is already standard in conventional cooling models. To go beyond this approximation, one would need to develop a full viscoelastic theory of matter incorporating heat conduction, in which the evolution equation for $T$ is coupled to those for $n$ and $u^{\mu}$. In that case, a simple heat-conduction equation such as (3) would no longer suffice to model the evolution.
@@ Acknowledgements
This work is partially supported by a Vanderbilt Seeding Success Grant, the IN2P3 Master Project NewMAC and the ANR Project GW-HNS (ANR-22-CE31-0001-01).